# New results related to cutters and to an extrapolated block-iterative method for finding a common fixed point of a collection of them

### v2

@@ 1.1. Background:
In this work we continue the study done in [22] by weakening one of the conditions made there about the relaxation parameters, namely we allow them to belong to a wider interval than the typically used interval $[\tau_{1},2-\tau_{2}]$, where $\tau_{1},\tau_{2}\in(0,1]$. In other words, we allow extrapolation. The appearance of extrapolation, while intended to accelerate the convergence of the algorithmic sequence by allowing deeper steps, introduces complications which do not appear without extrapolation, and it forces us to make a more restrictive assumption than in [22] on either the dynamic weights or on the common fixed point set in order to ensure global convergence. More precisely, we either require the weights to satisfy a certain intermittent condition (Condition 4.5 below) instead of imposing the condition $\sum_{k=0}^{\infty}w_{k}(i)=\infty$ for each $i\in I$ (Condition 4.4 below) as in [22] (following [1]), or we require the common fixed point set to have a nonempty interior (and then we allow Condition 4.4).
@@ 2. Preliminaries
[…]
@@
[…]
@@
(2.12) for all $\text{for all }q\in\textnormal{Fix}(T)\,\text{ and all }x\in X\,\text{ one has}\left\langle x-T(x),q-T(x)\right\rangle\leq 0.$
@@ 3. Results related to cutters
(3.3) $|||\mathbf{x}-\mathbf{P}_{\mathbf{Q}}(\mathbf{x})|||_{w}^{2}=\left(\min\{|||\mathbf{x}-\mathbf{z}|||_{w}\,|\,\mathbf{z}\in\mathbf{Q}\}\right)^{2}=\min\{|||\mathbf{x}-\mathbf{z}|||_{w}^{2}\,|\,\mathbf{z}\in\mathbf{Q}\}$
@@
(3.4) $|||\mathbf{x}-\mathbf{P}_{\mathbf{D}}(\mathbf{x})|||_{w}^{2}=\min\{|||\mathbf{x}-\mathbf{z}|||_{w}^{2}:\mathbf{z}\in\mathbf{D}\}=\min\{|||\mathbf{x}-\mathbf{J}(z)|||_{w}^{2}:z\in X\}\\ =\min\left\{\sum_{i\in I}w(i)\|x^{i}-z\|^{2}:z\in X\right\}.$
@@
Let $T_{1}$ and $T_{2}$ be two cutters defined on a real Hilbert space $Y$. If there exists some $x\in\textnormal{Fix}(T_{1}T_{2})\backslash\textnormal{Fix}(T_{2})$, then $\textnormal{Fix}(T_{1})\cap\textnormal{Fix}(T_{2})=\emptyset$.
@@
(3.7) $\mathbf{b}_{w}(\mathbf{x}):=\mathbf{x}+\frac{|||\mathbf{T}\left(\mathbf{x}\right)-\mathbf{x}|||_{w}^{2}}{|||\mathbf{P}_{\mathbf{D}}\left(\mathbf{T}\left(\mathbf{x}\right)\right)-\mathbf{x}|||_{w}^{2}}\left(\mathbf{P}_{\mathbf{D}}\left(\mathbf{T}\left(\mathbf{x}\right)\right)-\mathbf{x}\right),\quad\mathbf{x}\in\mathbf{D}\backslash\mathbf{F}$
@@
Note that $\mathbf{b}_{w}$ (from (3.7)) generalizes the expression $\mathbf{b}$ which appears in [46, Lemma 1.2] with orthogonal projections instead of general cutters we use here, and with constant weights ($w_{k}(i):=1/m$ for all $k\in\mathbb{N}\cup\{0\}$ and $i\in I$) and not general ones as we use here. In addition, and in Proposition 3.4 above, we filled a gap which appears in [46, Lemma 1.2] (see also [8, Theorem 2.8(i) and Corollary 2.12(i)] and [29, Proof of Proposition 2.4] for closely related results): the gap there is the unproven assertion that, with the notation we use here, $\mathbf{P_{D}}(\mathbf{P_{C}}(\mathbf{x}))\neq\mathbf{x}$ for all $\mathbf{x}\in\mathbf{D}\backslash\mathbf{C}$, where $\mathbf{C}:=C_{1}\times\cdots\times C_{m}$ and $C_{i}$ is a nonempty, closed and convex subset of $X$ for all $i\in I$; while it is claimed in [46, Lemma 1.2] that this assertion was proved in [46, Theorem 1.1(i)], as far as we understand, this is not the case. The reason that we indeed filled this gap follows from the observation that if we denote $T_{i}:=P_{C_{i}}$ for each $i\in I$ and consider the mapping $\mathbf{T}$ from (2.6), then Lemma 3.1 (with $Q_{i}:=C_{i}$, $i\in I$, and hence $\mathbf{Q}=\mathbf{C}$) guarantees that $\mathbf{T}=\mathbf{P_{C}}$, and since both $F_{i}:=\textnormal{Fix}(T_{i})=C_{i}$ for every $i\in I$ and $\mathbf{F}:=\textnormal{Fix}(\mathbf{T})=F_{1}\times\cdots\times F_{m}=\mathbf{C}$ (as follows from (2.7) and the equality $F_{i}=C_{i}$ for every $i\in I$), we conclude from Proposition 3.4 that $\mathbf{P}_{\mathbf{D}}\left(\mathbf{T}\left(\mathbf{x}\right)\right)\neq\mathbf{x}$, namely $\mathbf{P_{D}}(\mathbf{P_{C}}(\mathbf{x}))\neq\mathbf{x}$, as required.
@@
    (3.10) $\widehat{\lambda}_{x,w}:=\frac{|||\mathbf{T}\left(\mathbf{x}\right)-\mathbf{x}|||_{w}^{2}}{|||\mathbf{P}_{\mathbf{D}}\left(\mathbf{T}\left(\mathbf{x}\right)\right)-\mathbf{x}|||_{w}^{2}},$
@@
$=\frac{|||\mathbf{T}\left(\mathbf{x}\right)-\mathbf{x}|||_{w}^{2}}{|||\mathbf{P}_{\mathbf{D}}\left(\mathbf{T}\left(\mathbf{x}\right)\right)-\mathbf{x}|||_{w}^{2}}\Big{(}\left\langle\left\langle\mathbf{P}_{\mathbf{D}}\left(\mathbf{T}\left(\mathbf{x}\right)\right)-\mathbf{x,\mathbf{T}\left(\mathbf{x}\right)-P}_{\mathbf{D}}\left(\mathbf{T}\left(\mathbf{x}\right)\right)\right\rangle\right\rangle_{w}$
@@
(3.15) $=\frac{|||\mathbf{T}\left(\mathbf{x}\right)-\mathbf{x}|||_{w}^{2}}{|||\mathbf{P}_{\mathbf{D}}\left(\mathbf{T}\left(\mathbf{x}\right)\right)-\mathbf{x}|||_{w}^{2}}|||\mathbf{P}_{\mathbf{D}}\left(\mathbf{T}\left(\mathbf{x}\right)\right)-\mathbf{x}|||_{w}^{2}-|||\mathbf{T}\left(\mathbf{x}\right)-\mathbf{x}|||_{w}^{2}=0.$
@@
$|||\mathbf{P}_{\mathbf{D}}(\mathbf{T}(\mathbf{x}))-\mathbf{x}|||_{w}=|||\mathbf{P}_{\mathbf{D}}(\mathbf{T}(\mathbf{x}))-\mathbf{P}_{\mathbf{D}}(\mathbf{x})|||_{w}\leq|||\mathbf{T}(\mathbf{x})-\mathbf{x}|||_{w}.$
@@
$|||\mathbf{P}_{\mathbf{D}}(\mathbf{T}(\mathbf{x}))-\mathbf{x}|||_{w}^{2}=|||\mathbf{P}_{\mathbf{D}}(\mathbf{T}(\mathbf{x})-\mathbf{x})|||_{w}^{2}$
$=\left|\left|\left|\mathbf{J}\left(\sum_{i\in I}w(i)(T_{i}(x)-x)\right)\right|\right|\right|_{w}^{2}=\sum_{j\in I}w(j)\left\|\sum_{i\in I}w(i)(T_{i}(x)-x)\right\|^{2}$
@@
00 0 $\leq\left\langle\left\langle\mathbf{y}-\mathbf{T}\left(\mathbf{x}\right),\mathbf{T}\left(\mathbf{x}\right)-\mathbf{x}\right\rangle\right\rangle_{w}$
@@
$\left\langle\left\langle\mathbf{y}-\mathbf{b}_{w}(\mathbf{x)},\frac{|||\mathbf{P}_{\mathbf{D}}\left(\mathbf{T}\left(\mathbf{x}\right)\right)-\mathbf{x}|||_{w}^{2}}{|||\mathbf{T}\left(\mathbf{x}\right)-\mathbf{x}|||_{w}^{2}}\left(\mathbf{b}_{w}(\mathbf{x)-x}\right)\right\rangle\right\rangle_{w}\geq 0,\text{for all }\mathbf{y}\in\mathbf{F}\cap\mathbf{D},$
@@
00 0 $\leq\left\langle\left\langle\mathbf{y}-\mathbf{b}_{w}(\mathbf{x}),\mathbf{P}_{\mathbf{D}}(\mathbf{T}(\mathbf{x}))-\mathbf{x}\right\rangle\right\rangle_{w}$
@@
(3.20) $0\leq\epsilon|||\mathbf{x}-\mathbf{P}_{\mathbf{D}}(\mathbf{T}(\mathbf{x}))|||^{2}_{w}\leq\left\langle\left\langle\mathbf{y}-\mathbf{P}_{\mathbf{D}}(\mathbf{T}(\mathbf{x})),\mathbf{P}_{\mathbf{D}}(\mathbf{T}(\mathbf{x}))-\mathbf{x}\right\rangle\right\rangle_{w},$
@@
00 0 $\leq\langle u-T_{w,\widehat{\lambda}_{x,w}}(x),T_{w,\widehat{\lambda}_{x,w}}(x)-x\rangle=\left\langle u-T_{w,\widehat{\lambda}_{x,w}}(x),\widehat{\lambda}_{x,w}\left(T_{w}(x)-x\right)\right\rangle$
@@ 4. The block iterative extrapolated algorithm and the convergence theorems
Since $\tau_{1}/(2-\tau_{2})\leq 1/(2-1)\leq L(x^{k},w_{k})$ for all $k\in\mathbb{N}\cup\{0\}$ according to (2.2), it follows that for all $k\in\mathbb{N}\cup\{0\}$ the interval $[\tau_{1},(2-\tau_{2})L(x^{k},w_{k})]$ is nonempty, and hence one can indeed choose a sequence of relaxation parameters $\{\lambda_{k}\}_{k=0}^{\infty}$ in this interval. In standard BIP algorithms the parameters $\{\lambda_{k}\}_{k=0}^{\infty}$ are confined to the interval $\left[\tau_{1},2-\tau_{2}\right]$ for any user-chosen two real numbers $\tau_{1}$ and $\tau_{2}$ in the interval $(0,1],$ allowing under-relaxation or over-relaxation of the projections. Many papers investigate the role of relaxation parameters, see, e.g., [35]. Here, in Algorithm 4.1, the term $L(x^{k},w_{k})$ which might be greater than $1$, enables “deeper” steps which constitute the extrapolation beyond the usual interval that appears in convergence theorems for projection methods.
@@
(4.3) There are $\textnormal{There are }s\in\mathbb{N}\,\,\textnormal{and}\,\alpha>0\,\,\textnormal{such that both}\,\,\{I_{k}\}_{k=0}^{\infty}\,\,\textnormal{is}\,\,s\textnormal{-intermittent}\\ \textnormal{and}\,w_{k}(i)\geq\alpha\,\,\textnormal{for all}\,\,k\in\mathbb{N}\cup\{0\}\,\,\textnormal{and all}\,\,i\in I_{k},$
@@
Suppose that $m>1$. Let $s$ be an arbitrary even natural number and let $s^{\prime}:=s/2$. For each $t\in\mathbb{N}\cup\{0\}$ choose randomly, say using the uniform distribution on $\{ts^{\prime},ts^{\prime}+1,\ldots,ts^{\prime}+s^{\prime}-1\}$, a number $h_{t,1}\in\{ts^{\prime},ts^{\prime}+1,\ldots,ts^{\prime}+s^{\prime}-1\}$. Now choose (possibly randomly, using the uniform distribution on $[1/(2m),1/m]$) arbitrary real numbers $w_{h_{t,1}}(i)\in[1/(2m),1/m]$ for all $i\in I\backslash\{m\}$, and define $w_{h_{t,1}}(m):=1-\sum_{j=1}^{m-1}w_{h_{t,1}}(j)$. Now for each $k\in\{ts^{\prime},ts^{\prime}+1,\ldots,ts^{\prime}+s^{\prime}-1\}\backslash\{h_{t,1}\}$ and each $i\in I\backslash\{m\}$ choose an arbitrary (possibly randomly, using the uniform distribution on $[0,1/m]$) real number $w_{k}(i)\in[0,1/m]$, and define $w_{k}(m):=1-\sum_{j=1}^{m-1}w_{k}(j)$.
@@
$0<ts^{\prime}-\ell\leq k-\ell\leq ts+s^{\prime}-1-(t-1)s^{\prime}=2s^{\prime}-1=s-1,$
@@
Since $x_{\infty}$ is an accumulation point of $\{x^{k}\}_{k=0}^{\infty}$, there is a subsequence $\{x^{k_{t}}\}_{t=0}^{\infty}$ such that $x_{\infty}=\lim_{t\to\infty}x^{k_{t}}$. We need to show that $x_{\infty}\in F_{i}$ for all $i\in I$. Fix an arbitrary $i\in I$. By Condition 4.5, for all $t\in\mathbb{N}\cup\{0\}$ there is $h_{t,i}\in\{k_{t},k_{t}+1,\ldots,k_{t}+s-1\}$ such that $w_{h_{t,i}}(i)\geq\alpha$. Proposition 4.13 ensures that $\lim_{t\to\infty}\|x^{h_{t,i}}-x^{k_{t}}\|=0$, and hence $\|x^{h_{t,i}}-x_{\infty}\|\leq\|x^{h_{t,i}}-x^{k_{t}}\|+\|x^{k_{t}}-x_{\infty}\|\xrightarrow[t\to\infty]{}0+0=0$. Consider the sequence $\{(w_{h_{t,i}}(1),w_{h_{t,i}}(2),\ldots,w_{h_{t,i}}(m))\}_{t=0}^{\infty}$ of weight vectors in $[0,1]^{m}$. Since $[0,1]^{m}$ is a compact subset of $\mathbb{R}^{m}$, there is an infinite subset $N_{1}$ of $\mathbb{N}\cup\{0\}$ and a subsequence $\{(w_{h_{t,i}}(1),w_{h_{t,i}}(2),\ldots,w_{h_{t,i}}(m))\}_{t\in N_{1}}$ of $\{(w_{h_{t,i}}(1),w_{h_{t,i}}(2),\ldots,w_{h_{t,i}}(m))\}_{t=0}^{\infty}$ which converges to some weight vector $(w_{\infty,i}(1),w_{\infty,i}(2),\ldots,w_{\infty,i}(m))$. This vector satisfies $w_{\infty,i}(i)=\lim_{t\to\infty,t\in N_{1}}w_{h_{t,i}}(i)\geq\alpha$ because $w_{h_{t,i}}(i)\geq\alpha$ for all $t\in\mathbb{N}\cup\{0\}$ by the choice of $h_{t,i}$.

### v3

@@ 1.1. Background:
In this work we continue the study done in [22] by weakening one of the conditions made there about the relaxation parameters, namely we allow them to belong to a wider interval than the typically used interval $\tau_{1},2-\tau_{2}]$, where $\tau_{1},\tau_{2}\in(0,1]$. In other words, we allow extrapolation. The appearance of extrapolation, while intended to accelerate the convergence of the algorithmic sequence by allowing deeper steps, introduces complications which do not appear without extrapolation, and it forces us to make a more restrictive assumption than in [22] on either the dynamic weights or on the common fixed point set in order to ensure global convergence. More precisely, we either require the weights to satisfy a certain intermittent condition (Condition 4.5 below) instead of imposing the condition $\sum_{k=0}^{\infty}w_{k}(i)=\infty$ for each $i\in I$ (Condition 4.4 below) as in [22] (following [1]), or we require the common fixed point set to have a nonempty interior (and then we allow Condition 4.4).
@@ 2. Preliminaries
We use in the sequel the following notations and definitions. Unless otherwise stated, the ambient space is a real Hilbert space $X$ with an inner product $\langle\cdot,\cdot\rangle$ and an induced norm $\|\cdot\|$. Given $\rho>0$ and $x\in X$, we denote the closed ball with radius $\rho$ centered at $x$ by $B[x,\rho]:=\left\{y\in X\mid\left\|x-y\right\|\leq\rho\right\}$. Given a natural number $m$, define the index set $I$ as $I:=\left\{1,2,\ldots,m\right\}$. A function $w:I\rightarrow\left[0,1\right]$ which satisfies $\sum_{i\in I}w\left(i\right)=1$ is called a weight function, and if, in addition, $w(i)>0$ for all $i\in I$, then $w$ is called a positive weight function. We refer to the vector $(w(i))_{i\in I}\in\mathbb{R}^{m}$ as a weight vector, and when $w$ is a positive weight function, then we refer to $(w(i))_{i\in I}$ as a positive weight vector. Denote $\widehat{I}_{w}:=\{i\in I\mid w\left(i\right)>0\}$ and let $\widehat{w}:\widehat{I}_{w}\to(0,1]$ be the restriction of $w$ to $\widehat{I}_{w}$, that is, $\widehat{w}(i):=w(i)>0$ for all $i\in\widehat{I}_{w}$. Note that $\widehat{I}_{w}\neq\emptyset$ (since $\sum_{i\in I}w(i)=1$), that $\widehat{w}$ is a positive weight function defined on $\widehat{I}_{w}$, and the case $\widehat{w}(i)=1$ for some $i\in\widehat{I}$ is possible if and only if $\widehat{I}$ is a singleton.
@@
(2.9) $\text{Find a point }\mathbf{x}\text{ in }\mathbf{F}\cap\mathbf{D}\subseteq\mathbf{X}\mathbf{.}$
@@
[…]
@@ 3. Results related to cutters
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Note that $\mathbf{b}_{w}$ (from (3.7)) generalizes the expression $\mathbf{b}$ which appears in [46, Lemma 1.2] with orthogonal projections instead of general cutters as we use here, and with constant weights ($w_{k}(i):=1/m$ for all $k\in\mathbb{N}\cup\{0\}$ and $i\in I$) and not general ones as we use here. In addition, and in Proposition 3.4 above, we filled a gap which appears in [46, Lemma 1.2] (see also [8, Theorem 2.8(i) and Corollary 2.12(i)] and [29, Proof of Proposition 2.4] for closely related results): the gap there is the unproven assertion that, with the notation we use here, $\mathbf{P_{D}}(\mathbf{P_{C}}(\mathbf{x}))\neq\mathbf{x}$ for all $\mathbf{x}\in\mathbf{D}\backslash\mathbf{C}$, where $\mathbf{C}:=C_{1}\times\cdots\times C_{m}$ and $C_{i}$ is a nonempty, closed and convex subset of $X$ for all $i\in I$; while it is claimed in [46, Lemma 1.2] that this assertion was proved in [46, Theorem 1.1(i)], as far as we understand, this is not the case. The reason that we indeed filled this gap follows from the observation that if we denote $T_{i}:=P_{C_{i}}$ for each $i\in I$ and consider the mapping $\mathbf{T}$ from (2.6), then Lemma 3.1 (with $Q_{i}:=C_{i}$, $i\in I$, and hence $\mathbf{Q}=\mathbf{C}$) guarantees that $\mathbf{T}=\mathbf{P_{C}}$, and since both $F_{i}:=\textnormal{Fix}(T_{i})=C_{i}$ for every $i\in I$ and $\mathbf{F}:=\textnormal{Fix}(\mathbf{T})=F_{1}\times\cdots\times F_{m}=\mathbf{C}$ (as follows from (2.7) and the equality $F_{i}=C_{i}$ for every $i\in I$), we conclude from Proposition 3.4 that $\mathbf{P}_{\mathbf{D}}\left(\mathbf{T}\left(\mathbf{x}\right)\right)\neq\mathbf{x}$, namely $\mathbf{P_{D}}(\mathbf{P_{C}}(\mathbf{x}))\neq\mathbf{x}$, as required.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$0$ $\leq\left\langle\left\langle\mathbf{y}-\mathbf{T}\left(\mathbf{x}\right),\mathbf{T}\left(\mathbf{x}\right)-\mathbf{x}\right\rangle\right\rangle_{w}$
@@
[…]
@@
$0$ $\leq\left\langle\left\langle\mathbf{y}-\mathbf{b}_{w}(\mathbf{x}),\mathbf{P}_{\mathbf{D}}(\mathbf{T}(\mathbf{x}))-\mathbf{x}\right\rangle\right\rangle_{w}$
@@
[…]
@@
$0$ $\leq\langle u-T_{w,\widehat{\lambda}_{x,w}}(x),T_{w,\widehat{\lambda}_{x,w}}(x)-x\rangle=\left\langle u-T_{w,\widehat{\lambda}_{x,w}}(x),\widehat{\lambda}_{x,w}\left(T_{w}(x)-x\right)\right\rangle$
@@ 4. The block iterative extrapolated algorithm and the convergence theorems
Since $\tau_{1}/(2-\tau_{2})\leq 1/(2-1)\leq L(x^{k},w_{k})$ for all $k\in\mathbb{N}\cup\{0\}$ according to (2.2), it follows that for all $k\in\mathbb{N}\cup\{0\}$ the interval $\tau_{1},(2-\tau_{2})L(x^{k},w_{k})]$ is nonempty, and hence one can indeed choose a sequence of relaxation parameters $\{\lambda_{k}\}_{k=0}^{\infty}$ in this interval. In standard BIP algorithms the parameters $\{\lambda_{k}\}_{k=0}^{\infty}$ are confined to the interval $\left[\tau_{1},2-\tau_{2}\right]$ for any user-chosen two real numbers $\tau_{1}$ and $\tau_{2}$ in the interval $(0,1],$ allowing under-relaxation or over-relaxation of the projections. Many papers investigate the role of relaxation parameters, see, e.g., [35]. Here, in Algorithm 4.1, the term $L(x^{k},w_{k})$ which might be greater than $1$, enables “deeper” steps which constitute the extrapolation beyond the usual interval that appears in convergence theorems for projection methods.
@@
(4.3) $\textnormal{There are }s\in\mathbb{N}\,\,\textnormal{and}\,\alpha>0\,\,\textnormal{such that both}\,\,\{I_{k}\}_{k=0}^{\infty}\,\,\textnormal{is}\,\,s\textnormal{-intermittent}\\ \textnormal{and}\,w_{k}(i)\geq\alpha\,\,\textnormal{for all}\,\,k\in\mathbb{N}\cup\{0\}\,\,\textnormal{and all}\,\,i\in I_{k},$
@@
By doing this we obtain a sequence $\{w_{k}\}_{k=0}^{\infty}$ of weight functions which satisfies Condition 4.5 with $\alpha:=1/(2m)$ and $s$. Indeed, for every $\ell\in\mathbb{N}\cup\{0\}$ let $t:=\lceil\ell/s^{\prime}\rceil$ (where $\lceil\cdot\rceil$ is the strict ceiling function, which assigns to every $r\in\mathbb{R}$ the minimal integer which is strictly greater than $r$), and let $k:=h_{t,1}$. The definition of $h_{t,1}$ implies that $k\in\{ts^{\prime},ts^{\prime}+1,\ldots,ts^{\prime}+s^{\prime}-1\}$. In addition, the definition of $t$ implies that $(t-1)s^{\prime}\leq\ell<ts^{\prime}$. Hence,
$0<ts^{\prime}-\ell\leq k-\ell\leq ts^{\prime}+s^{\prime}-1-(t-1)s^{\prime}=2s^{\prime}-1=s-1,$
@@
Since $x_{\infty}$ is an accumulation point of $\{x^{k}\}_{k=0}^{\infty}$, there is a subsequence $\{x^{k_{t}}\}_{t=0}^{\infty}$ such that $x_{\infty}=\lim_{t\to\infty}x^{k_{t}}$. We need to show that $x_{\infty}\in F_{i}$ for all $i\in I$. Fix an arbitrary $i\in I$. By Condition 4.5, for all $t\in\mathbb{N}\cup\{0\}$ there is $h_{t,i}\in\{k_{t},k_{t}+1,\ldots,k_{t}+s-1\}$ such that $w_{h_{t,i}}(i)\geq\alpha$. Proposition 4.13 ensures that $\lim_{t\to\infty}\|x^{h_{t,i}}-x^{k_{t}}\|=0$, and hence $\|x^{h_{t,i}}-x_{\infty}\|\leq\|x^{h_{t,i}}-x^{k_{t}}\|+\|x^{k_{t}}-x_{\infty}\|\xrightarrow[t\to\infty]{}0+0=0$. Consider the sequence $\{(w_{h_{t,i}}(1),w_{h_{t,i}}(2),\ldots,w_{h_{t,i}}(m))\}_{t=0}^{\infty}$ of weight vectors in $0,1]^{m}$. Since $0,1]^{m}$ is a compact subset of $\mathbb{R}^{m}$, there is an infinite subset $N_{1}$ of $\mathbb{N}\cup\{0\}$ and a subsequence $\{(w_{h_{t,i}}(1),w_{h_{t,i}}(2),\ldots,w_{h_{t,i}}(m))\}_{t\in N_{1}}$ of $\{(w_{h_{t,i}}(1),w_{h_{t,i}}(2),\ldots,w_{h_{t,i}}(m))\}_{t=0}^{\infty}$ which converges to some weight vector $(w_{\infty,i}(1),w_{\infty,i}(2),\ldots,w_{\infty,i}(m))$. This vector satisfies $w_{\infty,i}(i)=\lim_{t\to\infty,t\in N_{1}}w_{h_{t,i}}(i)\geq\alpha$ because $w_{h_{t,i}}(i)\geq\alpha$ for all $t\in\mathbb{N}\cup\{0\}$ by the choice of $h_{t,i}$.